OpenStudy (anonymous):

Missing something simple! Check my work so far please? Diff EQs. The problem is \[\frac{ dy }{ dt } + \frac{ 2t }{ 1+t^2 }y = \frac{ 1 }{ 1+t^2 }\] So far, I move all terms to the left so: \[\frac{ dy }{ dt } + \frac{ 2t }{ 1+t^2 }y - \frac{ 1 }{ 1+t^2 } =0\] Next I divided both sides by y, giving : \[\frac{ \frac{ dy }{ dt } }{ y } + \frac{ 2t }{ 1+t^2 } - \frac{ 1 }{ y+yt^2 } = 0\] and now I am stuck

OpenStudy (anonymous):

I have the identity that \[\frac{ \frac{ dy }{ dt } }{ y } = \frac{ d }{ dt }\ln \left| y(t) \right|\] but I feel like I am missing a step before exponentiating each side

OpenStudy (anonymous):

Please, can anyone talk to me about differential equations?

OpenStudy (phi):

Do you know about Integrating Factors? a linear differential equation of the form \[ \frac{dy}{dt} +P(t) y = Q(t) \] has an integrating factor \[ e^{\int P(t) dt} \] the solution will be \[ e^{\int P(t) dt} y = \int e^{\int P(t) dt}Q(t) dt +c\]

OpenStudy (anonymous):

i just watched a video on them and had my mind blown. Thanks so much for the response, glad to know Im finally on the right track.

OpenStudy (anonymous):

So, I get \[1+x^2 \] as my integrating factor. My understanding is I multiply both sides by that now? So I would have \[\frac{ dy }{ dt }(1+t^2) + 2ty = 1\]

OpenStudy (anonymous):

Im not sure I understand what happens with the left side here, I think I did something wrong.

OpenStudy (phi):

I would start out as \[ \frac{ dy }{ dt } + \frac{ 2t }{ 1+t^2 }y = \frac{ 1 }{ 1+t^2 } \\ dy +\frac{ 2t }{ 1+t^2 }y \ dt= \frac{ 1 }{ 1+t^2 } \ dt \] now multiply by the integrating factor \[ (1+t^2) dy +(1+t^2)\frac{ 2t }{ 1+t^2 }y \ dt= (1+t^2)\frac{ 1 }{ 1+t^2 } \ dt \\ (1+t^2) dy + 2t\ y\ dt = dt \] now integrate both sides. the left side is d (f(t) y) = f(t) dy + y d(f(t))

OpenStudy (phi):

or if we pick up where you left off \[ \frac{ dy }{ dt }(1+t^2) + 2ty = 1 \\ (1+t^2) dy + 2 t \ y\ dt = dt \]

OpenStudy (anonymous):

so multiply both sides by dt?

OpenStudy (phi):

yes, we almost always split the differentials

OpenStudy (anonymous):

ahhhh ok, I was wondering but that makes sense, back to the basics

OpenStudy (phi):

though it may not be obvious the left side is \[ d( (1+t^2) y) \] and you have \[ d( (1+t^2) y)= dt \\ \int d( (1+t^2) y)= \int dt \]

OpenStudy (anonymous):

what happens to the 2tydt on the left side?

OpenStudy (phi):

the product rule (take the derivative of the product of two variables) is d (u v) = u dv + v du in this problem we go in the other direction u dv + v du = d (u v)

OpenStudy (phi):

in other words if you expand d ( (1+t^2) y) you get (1+t^2) dy + y 2t dt

OpenStudy (anonymous):

arg ok, I think I get that. I'll play around with them until it sinks in.

OpenStudy (phi):

\[ \int d( (1+t^2) y)= \int dt \\ (1+t^2) y = t+C \\ y= \frac{t}{1+t^2}+C' \] where we can rename C/(1+t^2) as a new constant of integration C'